# Large Language Models in the Task of Automatic Validation of Text Classifier Predictions

### v1

@@ Abstract
Machine learning models for text classification are trained to predict a class for a given text. To do this, training and validation samples must be prepared: a set of texts is collected, and each text is assigned a class. These classes are usually assigned by human annotators with different expertise levels, depending on the specific classification task. Collecting such samples from scratch is labor-intensive because it requires finding specialists and compensating them for their work; moreover, the number of available specialists is limited, and their productivity is constrained by human factors. While it may not be too resource-intensive to collect samples once, the ongoing need to retrain models (especially in incremental learning pipelines [incremental1]) to address data drift (also called model drift [ibm_model_drift]) makes the data collection process crucial and costly over the model’s entire lifecycle. This paper proposes several approaches to replace human annotators with Large Language Models (LLMs) to test classifier predictions for correctness, helping ensure model quality and support high-quality incremental learning.
@@ 1 Introduction
With the growing popularity of LLMs, more professionals are now exploring whether humans can be replaced in certain areas of work. Because LLMs possess knowledge and reasoning abilities in many domains and can handle free-format tasks [awesome_prompt1][awesome_prompt2], they are strong candidates to augment or even replace human annotation. Simply replacing humans with another text classifier (to validate predictions of the first classifier) would lead to a conventional classification pipeline, requiring frequent retraining of two models.
@@ 2 Related Work
In research, opportunities for using LLMs in classification tasks can be categorized into two groups: Large Language Models as Classifiers (In-Context Learning)[icl] and LLM as Data Annotators.
@@ 2.1 LLM ICL Classifier Replacement Approaches
"In-context learning (ICL) classifier replacement"refers to using LLMs to directly perform classification by prompting the model with examples and letting it predicts class labels, instead of training a separate classifier. Recent studies have explored using powerful LLMs in a few-shot or zero-shot manner to replace traditional classifiers. In the paper [breakingbankchatgptfewshot], the authors compared the LLM with the ICL approach and fine-tuned masked language models on 77-class classification problem (on the banking77 dataset [banking77]). The authors showed that GPT-4 [gpt4], using a few well-selected examples and a limited number of examples per class, outperformed the considered fine-tuned models, opening the way for building rapid solutions under conditions of limited computational resources and data. One major drawback is the cost of using proprietary models. Although proprietary models become cheaper annually (the authors cited around  \$1,600 for intensive GPT-4 usage), their use can still be expensive for large companies.
The authors of the paper [yu2023openclosedsmalllanguage] compared LLMs with fine-tuned approaches of classical language models on Named Entity Recognition (NER) and ordinary classification tasks. The authors found that closed-source LLMs (GPT-4 in particular) achieved the highest performance on challenging, "generalization-heavy"tasks, but smaller fine-tuned models often matched or surpassed LLMs on easier datasets. Additionally, pre-trained open-source LLMs (Llama-2 [llama2]) occasionally yielded better results than GPT-3.5 [gpt35].
In a relevant study [schnabel2025multistagelargelanguagemodel], the authors explored the potential of replacing humans in the task of classifying query-to-document relationships (matching relevant documents to queries). Their proposed multi-stage approach using GPT-4o [gpt4o] demonstrated that smart prompting and sequential queries with varied prompts could enhance LLM performance without pre-training methods.
@@ 2.2 LLM as Data Annotator
For instance, authors in [Gilardi_2023] examined replacing human annotators with GPT-3.5 in a zero-shot setting, achieving higher annotation quality than humans in 4 out of 5 task types, increased consistency with carefully collected human annotations, and approximately a 20-fold cost reduction compared to hired annotators. The paper [pangakis2024knowledgedistillationautomatedannotation] utilized LLM-generated annotations (termed distillation) to label training data, confirming that LLM-based labels offer "a fast, efficient, and cost-effective method" nearly equaling human-supervised models.
The authors of [ALMOHAMAD20252402] clearly demonstrated that LLM annotation can successfully classify image data (X-rays), resulting in high-quality CNN models. The paper [rouzegar2024enhancingtextclassificationllmdriven] proposed using GPT-3.5 to selectively replace human annotation only when confident, creating confidence thresholds for each dataset. However, the method relies on GPT-3.5 self-estimating its confidence, which can be unreliable [li2024thinktwicetrustingselfdetection].
@@
  • Many studies use data that may have already been used in the training samples of most LLMs before the papers were written. For example, datasets used in [breakingbankchatgptfewshot], [schnabel2025multistagelargelanguagemodel], and [rouzegar2024enhancingtextclassificationllmdriven] are publicly available and likely part of LLM training. The data in this paper is unique and described in Section 3.1.
@@ 3.4.2 Annotation metrics
In the annotation task, the only effective way to assess an LLM annotation capability is to measure how often its predictions coincide with those of human annotators, since the models initially have access only to human‐provided labels. Confidence in human annotation can be increased by collecting labels from several annotators on the same instance and retaining it only if they all agree, thus forming a high‐quality test set. Let $N$ be the total number of samples, $y_{i}$ be the prediction of annotator ${H}$ (human) or ${M}$ (LLM). Since human annotation on the benchmark was collected only twice (data collection from experts and data collection from regular annotators), we will fix $H$.
@@ 3.6.1 Zero-Shot and Few-Shot
LLM is capable of solving problems without any examples [kojima2023largelanguagemodelszeroshot], only based on the knowledge learned during training. But as practice of application and many researches show, if you give some examples of problem solving [fewshot], the quality of solved problems and following the instructions increases. Positive examples are defined for each intent, negative examples are generated specifically for this work as examples that are most difficult to classify by models and humans.
@@ 3.7 Alignment (SFT)
Two approaches were used to improve annotation quality: soft fine‑tuning on the classification task (adding a classification layer directly to the LLM) and soft fine‑tuning to enhance the LLM reasoning ability for correct annotation. To avoid modifying all LLM parameters due to computational resource constraints, we employ a low‑rank adaptation (LoRA) [lora]
@@ 3.8.1 Design
RAG typically comprises two components: retrieval and re‑ranking. Retrieval selects a set of documents based on simple metrics (e.g., cosine similarity between encoder embeddings of the query and the documents). Although retrieval is fast, it does not order documents precisely by relevance. Retrieval can also be performed using statistical algorithms without neural networks (e.g., BM25 [bm25]). Retrieved documents are usually stored in a vector database. These documents can then be re‑ranked by a more complex model (cross‑encoders) which helps to select the most relevant documents and order them by importance to the query. Large documents are split into smaller ones using various techniques (by paragraph, special symbols, or sentence-similarity within a chunk). This chunking ensures each piece is small enough to include in full without further division.
@@ 3.11 Annotation completion time
Previous experiments conducted by the company showed that 1 multi-class annotation takes 13.4 seconds on average, while 1 binary annotation takes 6.8 seconds on average (these statistics take into account that several annotators can annotate tasks in parallel). When using vLLM [vllm] and 1 GPU per 1 multi-class, the best approach with the ensemble takes  0.85 seconds, which is 8 times faster than the binary annotation task and 15.7 times faster than multi-class annotation.
@@ 5.1 LLMs
  • Vikhr NeMo-12b [vikhr]
    Bilingual Russian-English instruction Mistral-Nemo[mistral] with a 40k SentencePiece vocab adapted to Russian.
@@ 5.2.1 Retrievals
  • multilingual-e5-large-instruct [e5]
@@ 5.2.3 Embedding storage
Document embeddings are stored in the faiss IndexFlatIP [faiss]. This index stores all vectors in a "flat"form, without additional structure or compression, and when searching, simply calculates the scalar product between the query and each vector in the index. This approach guarantees accurate search for the closest neighbors according to the scalar product metric.
@@ 5.3.2 SFT (reasonings)
The collected training data using GPT-4o [gpt4o] and Deepseek-R1[deepseek] were used to SFT the Qwen2.5-32b model. The reasoning of other Large Language Models was parsed further, with the reasoning placed in the <think></think> block and the answer placed in <answer></answer>. Generations were only used for training if they led to the correct answer (the LLM answer matched the answer of professional annotators with extensive experience). it is also important not only to be able to reason well about the question, but also to be able to work with the information given in the task. Therefore, documents are also submitted in the prompt using RAG (Section 3.8) in the <retrieved></retrieved> block. It is important to note that the training data does not overlap with the benchmark data.
@@ 6.2.1 Multi-class
Model Coverage (conf="unk") Coverage ($\neq$"unk") Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 6.3 Prob approach after reasonings
The initial distribution (Fig. 6.3) is multimodal in the form of $U$ with peaks at 0 and 1. LLM is generally confident in its answers and separates them well into two distinct buckets. When changing the approach, it can be seen that the examples in which the model is less confident become fewer because they diverge towards the peaks 0 and 1, and the answers become more deterministic. This is especially true for values in the range 1 -- $\epsilon$, where the threshold is selected, because the maximum probability among the other answers is taken, and the answers "yes"and "no"are opposite in meaning, and LLM tends to give the maximum probability to one or the another answer. Finally, the change in distribution means that the model is highly confident in its answer, since all the necessary reasoning for making a decision was provided, even if it was incorrect (model hallucinations); therefore, using classification thresholds for a probabilistic approach is no longer effective.
@@ 6.4.1 Multi-class
Model Coverage (conf="unk") Coverage ($\neq$"unk") Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@
Model Coverage (conf="unk") Coverage ($\neq$"unk") Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 6.5 SFT (classification)
Model Coverage (conf="unk") Coverage ($\neq$"unk") Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 6.7.1 Multi-class
Model Coverage (conf="unk") Coverage ($\neq$"unk") Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 8 Conclusion
[…]
[heading=bibintoc]

### v2

@@ Abstract
Machine learning models for text classification are trained to predict a class for a given text. To do this, training and validation samples must be prepared: a set of texts is collected, and each text is assigned a class. These classes are usually assigned by human annotators with different expertise levels, depending on the specific classification task. Collecting such samples from scratch is labor-intensive because it requires finding specialists and compensating them for their work; moreover, the number of available specialists is limited, and their productivity is constrained by human factors. While it may not be too resource-intensive to collect samples once, the ongoing need to retrain models (especially in incremental learning pipelines [1]) to address data drift (also called model drift [2]) makes the data collection process crucial and costly over the model’s entire lifecycle. This paper proposes several approaches to replace human annotators with Large Language Models (LLMs) to test classifier predictions for correctness, helping ensure model quality and support high-quality incremental learning.
@@ 1 Introduction
With the growing popularity of LLMs, more professionals are now exploring whether humans can be replaced in certain areas of work. Because LLMs possess knowledge and reasoning abilities in many domains and can handle free-format tasks [3][4], they are strong candidates to augment or even replace human annotation. Simply replacing humans with another text classifier (to validate predictions of the first classifier) would lead to a conventional classification pipeline, requiring frequent retraining of two models.
@@ 2 Related Work
In research, opportunities for using LLMs in classification tasks can be categorized into two groups: Large Language Models as Classifiers (In-Context Learning)[5] and LLM as Data Annotators.
@@ 2.1 LLM ICL Classifier Replacement Approaches
"In-context learning (ICL) classifier replacement"refers to using LLMs to directly perform classification by prompting the model with examples and letting it predicts class labels, instead of training a separate classifier. Recent studies have explored using powerful LLMs in a few-shot or zero-shot manner to replace traditional classifiers. In the paper [6], the authors compared the LLM with the ICL approach and fine-tuned masked language models on 77-class classification problem (on the banking77 dataset [7]). The authors showed that GPT-4 [8], using a few well-selected examples and a limited number of examples per class, outperformed the considered fine-tuned models, opening the way for building rapid solutions under conditions of limited computational resources and data. One major drawback is the cost of using proprietary models. Although proprietary models become cheaper annually (the authors cited around  \$1,600 for intensive GPT-4 usage), their use can still be expensive for large companies.
The authors of the paper [9] compared LLMs with fine-tuned approaches of classical language models on Named Entity Recognition (NER) and ordinary classification tasks. The authors found that closed-source LLMs (GPT-4 in particular) achieved the highest performance on challenging, "generalization-heavy"tasks, but smaller fine-tuned models often matched or surpassed LLMs on easier datasets. Additionally, pre-trained open-source LLMs (Llama-2 [10]) occasionally yielded better results than GPT-3.5 [11].
In a relevant study [12], the authors explored the potential of replacing humans in the task of classifying query-to-document relationships (matching relevant documents to queries). Their proposed multi-stage approach using GPT-4o [13] demonstrated that smart prompting and sequential queries with varied prompts could enhance LLM performance without pre-training methods.
@@ 2.2 LLM as Data Annotator
For instance, authors in [14] examined replacing human annotators with GPT-3.5 in a zero-shot setting, achieving higher annotation quality than humans in 4 out of 5 task types, increased consistency with carefully collected human annotations, and approximately a 20-fold cost reduction compared to hired annotators. The paper [15] utilized LLM-generated annotations (termed distillation) to label training data, confirming that LLM-based labels offer "a fast, efficient, and cost-effective method" nearly equaling human-supervised models.
The authors of [16] clearly demonstrated that LLM annotation can successfully classify image data (X-rays), resulting in high-quality CNN models. The paper [17] proposed using GPT-3.5 to selectively replace human annotation only when confident, creating confidence thresholds for each dataset. However, the method relies on GPT-3.5 self-estimating its confidence, which can be unreliable [18].
@@
  • Many studies use data that may have already been used in the training samples of most LLMs before the papers were written. For example, datasets used in [6], [12], and [17] are publicly available and likely part of LLM training. The data in this paper is unique and described in Section 3.1.
@@ 3.4.2 Annotation metrics
In the annotation task, the only effective way to assess an LLM annotation capability is to measure how often its predictions coincide with those of human annotators, since the models initially have access only to human‐provided labels. Confidence in human annotation can be increased by collecting labels from several annotators on the same instance and retaining it only if they all agree, thus forming a high‐quality test set. Let $N$ be the total number of samples, $y_{i}$ be the prediction of annotator H{H} (human) or M{M} (LLM). Since human annotation on the benchmark was collected only twice (data collection from experts and data collection from regular annotators), we will fix $H$.
@@ 3.6.1 Zero-Shot and Few-Shot
LLM is capable of solving problems without any examples [19], only based on the knowledge learned during training. But as practice of application and many researches show, if you give some examples of problem solving [20], the quality of solved problems and following the instructions increases. Positive examples are defined for each intent, negative examples are generated specifically for this work as examples that are most difficult to classify by models and humans.
@@ 3.7 Alignment (SFT)
Two approaches were used to improve annotation quality: soft fine‑tuning on the classification task (adding a classification layer directly to the LLM) and soft fine‑tuning to enhance the LLM reasoning ability for correct annotation. To avoid modifying all LLM parameters due to computational resource constraints, we employ a low‑rank adaptation (LoRA) [21]
@@ 3.8.1 Design
RAG typically comprises two components: retrieval and re‑ranking. Retrieval selects a set of documents based on simple metrics (e.g., cosine similarity between encoder embeddings of the query and the documents). Although retrieval is fast, it does not order documents precisely by relevance. Retrieval can also be performed using statistical algorithms without neural networks (e.g., BM25 [22]). Retrieved documents are usually stored in a vector database. These documents can then be re‑ranked by a more complex model (cross‑encoders) which helps to select the most relevant documents and order them by importance to the query. Large documents are split into smaller ones using various techniques (by paragraph, special symbols, or sentence-similarity within a chunk). This chunking ensures each piece is small enough to include in full without further division.
@@ 3.11 Annotation completion time
Previous experiments conducted by the company showed that 1 multi-class annotation takes 13.4 seconds on average, while 1 binary annotation takes 6.8 seconds on average (these statistics take into account that several annotators can annotate tasks in parallel). When using vLLM [23] and 1 GPU per 1 multi-class, the best approach with the ensemble takes  0.85 seconds, which is 8 times faster than the binary annotation task and 15.7 times faster than multi-class annotation.
@@ 5.1 LLMs
  • Vikhr NeMo-12b [26]
    Bilingual Russian-English instruction Mistral-Nemo[27] with a 40k SentencePiece vocab adapted to Russian.
@@ 5.2.1 Retrievals
  • multilingual-e5-large-instruct [30]
@@ 5.2.3 Embedding storage
Document embeddings are stored in the faiss IndexFlatIP [33]. This index stores all vectors in a "flat"form, without additional structure or compression, and when searching, simply calculates the scalar product between the query and each vector in the index. This approach guarantees accurate search for the closest neighbors according to the scalar product metric.
@@ 5.3.2 SFT (reasonings)
The collected training data using GPT-4o [13] and Deepseek-R1[34] were used to SFT the Qwen2.5-32b model. The reasoning of other Large Language Models was parsed further, with the reasoning placed in the <think></think> block and the answer placed in <answer></answer>. Generations were only used for training if they led to the correct answer (the LLM answer matched the answer of professional annotators with extensive experience). it is also important not only to be able to reason well about the question, but also to be able to work with the information given in the task. Therefore, documents are also submitted in the prompt using RAG (Section 3.8) in the <retrieved></retrieved> block. It is important to note that the training data does not overlap with the benchmark data.
@@ 6.2.1 Multi-class
Model $Coverage (conf\neq"unk")$ Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 6.3 Prob approach after reasonings
The initial distribution (Fig. 6.3) is multimodal in the form of $U$ with peaks at 0 and 1. LLM is generally confident in its answers and separates them well into two distinct buckets. When changing the approach, it can be seen that the examples in which the model is less confident become fewer because they diverge towards the peaks 0 and 1, and the answers become more deterministic. This is especially true for values in the range 1 - $\epsilon$, where the threshold is selected, because the maximum probability among the other answers is taken, and the answers "yes"and "no"are opposite in meaning, and LLM tends to give the maximum probability to one or the another answer. Finally, the change in distribution means that the model is highly confident in its answer, since all the necessary reasoning for making a decision was provided, even if it was incorrect (model hallucinations); therefore, using classification thresholds for a probabilistic approach is no longer effective.
@@ 6.4.1 Multi-class
Model $Coverage (conf\neq"unk")$ Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@
Model $Coverage (conf\neq"unk")$ Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 6.5 SFT (classification)
Model $Coverage (conf\neq"unk")$ Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 6.7.1 Multi-class
Model $Coverage (conf\neq"unk")$ Accuracy (w/o conf) Accuracy (conf=wrong) $\text{Accuracy}_{\mathrm{total}}$
@@ 8 Conclusion
Список литературы (или источников)
  • [1] Jiangpeng He, Runyu Mao, Zeman Shao and Fengqing Zhu ‘‘Incremental Learning In Online Scenario’’, 2021 arXiv: https://arxiv.org/abs/2003.13191
  • [2] IBM ‘‘Model Drift: Why It Happens and How to Detect It’’ Accessed: 2024-07-16 URL: https://www.ibm.com/think/topics/model-drift
  • [3] Pranab Sahoo, Ayush Kumar Singh, Sriparna Saha, Vinija Jain, Samrat Mondal and Aman Chadha ‘‘A Systematic Survey of Prompt Engineering in Large Language Models: Techniques and Applications’’, 2025 arXiv: https://arxiv.org/abs/2402.07927
  • [4] Shubham Vatsal and Harsh Dubey ‘‘A Survey of Prompt Engineering Methods in Large Language Models for Different NLP Tasks’’, 2024 arXiv: https://arxiv.org/abs/2407.12994
  • [5] Tom B Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever and Dario Amodei ‘‘Language Models are Few-Shot Learners’’, 2020 eprint: 2005.14165
  • [6] Lefteris Loukas, Ilias Stogiannidis, Prodromos Malakasiotis and Stavros Vassos ‘‘Breaking the Bank with ChatGPT: Few-Shot Text Classification for Finance’’, 2023 arXiv: https://arxiv.org/abs/2308.14634
  • [7] Iñigo Casanueva, Tadas Temčinas, Daniela Gerz, Matthew Henderson and Ivan Vulić ‘‘Efficient Intent Detection with Dual Sentence Encoders’’, 2020 arXiv: https://arxiv.org/abs/2003.04807
  • [8] OpenAI ‘‘GPT-4 Technical Report’’ In CoRR abs/2303.08774, 2023 DOI: 10.48550/ARXIV.2303.08774
  • [9] Hao Yu, Zachary Yang, Kellin Pelrine, Jean Francois Godbout and Reihaneh Rabbany ‘‘Open, Closed, or Small Language Models for Text Classification?’’, 2023 arXiv: https://arxiv.org/abs/2308.10092
  • [10] Hugo Touvron, Louis Martin, Kevin Stone, Paul Albert, Abdullah Almahairi, Yacine Babaei, Nikita Bashlykov, Suchin Batra, Priya Bhargava, Shagun Bhosale, … and Thomas Scialom ‘‘Llama 2: Open Foundation and Fine‑Tuned Chat Models’’ In CoRR abs/2307.09288, 2023 DOI: 10.48550/ARXIV.2307.09288
  • [11] Junjie Ye, Xuanting Chen, Nuo Xu, Can Zu, Zekai Shao, Yuhan Liu, Zeyang Zhou, Chao Gong, Jie Shen, Siming Chen, Tao Gui, Qi Zhang and Xuanjing Huang ‘‘A Comprehensive Capability Analysis of GPT‑3 and GPT‑3.5 Series Models’’ In CoRR abs/2303.10420, 2023 DOI: 10.48550/ARXIV.2303.10420
  • [12] Julian A. Schnabel, Johanne R. Trippas, Falk Scholer and Danula Hettiachchi ‘‘Multi-stage Large Language Model Pipelines Can Outperform GPT-4o in Relevance Assessment’’, 2025 arXiv: https://arxiv.org/abs/2501.14296
  • [13] OpenAI ‘‘GPT-4o System Card’’ Accessed: 2025-05-16, https://openai.com/index/gpt-4o-system-card/, 2024
  • [14] Fabrizio Gilardi, Meysam Alizadeh and Maël Kubli ‘‘ChatGPT outperforms crowd workers for text-annotation tasks’’ In Proceedings of the National Academy of Sciences 120.30 Proceedings of the National Academy of Sciences, 2023 DOI: 10.1073/pnas.2305016120
  • [15] Nicholas Pangakis and Samuel Wolken ‘‘Knowledge Distillation in Automated Annotation: Supervised Text Classification with LLM-Generated Training Labels’’, 2024 arXiv: https://arxiv.org/abs/2406.17633
  • [16] Fares Al Mohamad, Leonhard Donle, Felix Dorfner, Laura Romanescu, Kristin Drechsler, Mike P. Wattjes, Jawed Nawabi, Marcus R. Makowski, Hartmut Häntze, Lisa Adams, Lina Xu, Felix Busch, Aymen Meddeb and Keno Kyrill Bressem ‘‘Open-source Large Language Models can Generate Labels from Radiology Reports for Training Convolutional Neural Networks’’ In Academic Radiology 32.5, 2025, pp. 2402–2410 DOI: https://doi.org/10.1016/j.acra.2024.12.028
  • [17] Hamidreza Rouzegar and Masoud Makrehchi ‘‘Enhancing Text Classification through LLM-Driven Active Learning and Human Annotation’’, 2024 arXiv: https://arxiv.org/abs/2406.12114
  • [18] Moxin Li, Wenjie Wang, Fuli Feng, Fengbin Zhu, Qifan Wang and Tat-Seng Chua ‘‘Think Twice Before Trusting: Self-Detection for Large Language Models through Comprehensive Answer Reflection’’, 2024 arXiv: https://arxiv.org/abs/2403.09972
  • [19] Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo and Yusuke Iwasawa ‘‘Large Language Models are Zero-Shot Reasoners’’, 2023 arXiv: https://arxiv.org/abs/2205.11916
  • [20] Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever and Dario Amodei ‘‘Language Models are Few-Shot Learners’’, 2020 arXiv: https://arxiv.org/abs/2005.14165
  • [21] Edward J Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang and Weizhu Chen ‘‘LoRA: Low-Rank Adaptation of large language models’’, 2021 eprint: 2106.09685
  • [22] Stephen E. Robertson and Hugo Zaragoza ‘‘The probabilistic relevance framework: BM25 and beyond’’ In Foundations and Trends in Information Retrieval 3.4, 2009, pp. 333–389
  • [23] Woosuk Kwon, Zhuohan Li, Siyuan Zhuang, Ying Sheng, Lianmin Zheng, Cody Hao Yu, Joseph E. Gonzalez, Hao Zhang and Ion Stoica ‘‘Efficient Memory Management for Large Language Model Serving with PagedAttention’’ In Proceedings of the ACM SIGOPS 29th Symposium on Operating Systems Principles, 2023
  • [24] Laurens Van Der Maaten ‘‘The Llama 3 Herd of Models’’ In arXiv preprint arXiv:2407.21783, 2024
  • [25] Qwen Team ‘‘Qwen2.5: A Party of Foundation Models’’, 2024 URL: https://qwenlm.github.io/blog/qwen2.5/
  • [26] Aleksandr Nikolich, Konstantin Korolev, Sergei Bratchikov, Igor Kiselev and Artem Shelmanov ‘‘Vikhr: Constructing a State-of-the-art Bilingual Open-Source Instruction-Following Large Language Model for Russian’’ In Proceedings of the 4rd Workshop on Multilingual Representation Learning (MRL) @ EMNLP-2024 Association for Computational Linguistics, 2024 URL: https://arxiv.org/pdf/2405.13929
  • [27] NVIDIA Corporation and Mistral AI ‘‘NeMo Mistral Toolkit for Large Language Models’’ Accessed: 2025-05-16, https://developer.nvidia.com/nemo/mistral, 2024
  • [28] Gemma Team ‘‘Gemma’’ Kaggle, 2024 DOI: 10.34740/KAGGLE/M/3301
  • [29] Nexusflow ‘‘Introducing Athene‑V2: Advancing Beyond the Limits of Scaling with Targeted Post‑training’’ Blog post announcing Athene‑V2‑Chat‑72B, https://nexusflow.ai/blogs/athene-v2, 2024
  • [30] Liang Wang, Nan Yang, Xiaolong Huang, Linjun Yang, Rangan Majumder and Furu Wei ‘‘Multilingual E5 Text Embeddings: A Technical Report’’ In arXiv preprint arXiv:2402.05672, 2024
  • [31] Jianlv Chen, Shitao Xiao, Peitian Zhang, Kun Luo, Defu Lian and Zheng Liu ‘‘BGE M3-Embedding: Multi-Lingual, Multi-Functionality, Multi-Granularity Text Embeddings Through Self-Knowledge Distillation’’, 2024 arXiv:2402.03216 [cs.CL]
  • [32] Artem Snegirev, Maria Tikhonova, Anna Maksimova, Alena Fenogenova and Alexander Abramov ‘‘The Russian-focused embedders’ exploration: ruMTEB benchmark and Russian embedding model design’’, 2024 arXiv:2408.12503 [cs.CL]
  • [33] Matthijs Douze, Alexandr Guzhva, Chengqi Deng, Jeff Johnson, Gergely Szilvasy, Pierre-Emmanuel Mazaré, Maria Lomeli, Lucas Hosseini and Hervé Jégou ‘‘The Faiss library’’, 2024 arXiv:2401.08281 [cs.LG]
  • [34] DeepSeek-AI et al. ‘‘DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning’’, 2025 arXiv: https://arxiv.org/abs/2501.12948